Use the vertex
(h, k) (-4,-1)
and a point on the graph
(x, y) (-6,3)
to find the general form of the equation of the quadratic function.

Answer by ikleyn(52781)

.
From the condition, using the info about the vertex, you can write


y = a*(x-(-4))^2 - 1,    or,   which is the same


y = a*(x+4)^2 -1,


where "a" is unknown real coefficient.  To find "a", use given condition on another point. It gives you an equation

3 = a*((-6)+4)^2 -1,  which is equivalent to


4 = a*2^2,   or   4 = 4a.


Hence,  a= 1,   and your quadratic function is   y = (x+4))^2 - 1.

Solved.